please help me with this problem:
Multiply (a+b)(a^2+ab+b^2)

Answer by Alan3354(69443)

please help me with this problem:
Multiply (a+b)(a^2+ab+b^2)
---------------
= a*(a^2+ab+b^2) + b*(a^2+ab+b^2)
= a^3 = a^2b + ab^2 + a^2b + ab^2 + b^3
collect terms
= a^3 + 2a^2b + 2ab^2 + b^3